OpenStudy (anonymous):

rewrite Y=√16X-32(all of that square rooted) -3 to make it easy to graph using translation.

OpenStudy (anonymous):

translations 2 left and 3 down translations 2 Right and 3 Down and the rewritten may be 4√X-2, y=√X-2 Y=√X+2

OpenStudy (anonymous):

those are possible answers for the two parts